1)A rod of length(, located along the y-axis has a total charge Q and a uniform linear charge density (. Find the electric potential at a point located on the x-axis a distance x from the origin.



1
Expert's answer
2021-07-25T09:22:59-0400

"x-axis distance\\l=x\\Charge=Q\\\lambda=\frac{q}{l}"

Electric potential

"V=\frac{kq}{r}=\frac{k\times\lambda l}{x}" Where l=x

"V=\frac{k\lambda x}{x}"

V=k"\lambda"

"V=\frac{1}{4\pi\epsilon_0}\lambda"